OpenStudy (anonymous):

Let g(x)=6x-x^2+14sqrt(x)-15. Find the slope of the tangent line to the graph of g(x) at x=4.

OpenStudy (jdoe0001):

\(\bf g(x)=6x-x^2+14\sqrt{x}-15\qquad \textit{use the power rule}\\ \quad \\ \textit{keep in mind that}\quad \sqrt{x}\implies \large x^{\frac{1}{2}}\)

OpenStudy (anonymous):

I'll see if I can find that in my notes... thanks.

OpenStudy (anonymous):

But what do I do with the g(x)=4?

OpenStudy (jdoe0001):

once you find the derivative, you replace x = 4 :) g(x) = f'(x)

OpenStudy (anonymous):

If I do the power rule I get... 6-2x+7x^(-1/2). Is that right? And what do I do from there if it is?

OpenStudy (jdoe0001):

welp... that came out off... well.. once you have the derivative , just set x = 4, f'(4)

OpenStudy (anonymous):

I got it wrong? That's not the derivative?

OpenStudy (jdoe0001):

you're correct, that's what I got

OpenStudy (anonymous):

What did I do wrong? 6x -> 6. -x^2 -> -2x. 14x^(1/2) -> 7x^-1/2, no?

OpenStudy (jdoe0001):

\(\bf g(x)=6x-x^2+14\sqrt{x}-15\\ \quad \ g'(x) = 6-2x+7x^{\frac{1}{2}}\qquad then\quad g'(4)\)

OpenStudy (jdoe0001):

hmm forgot the -

OpenStudy (anonymous):

Oh. So now just plug in 4? And thats the answer?

OpenStudy (jdoe0001):

\(\bf g(x)=6x-x^2+14\sqrt{x}-15\\ \quad \\ g'(x) = 6-2x+7x^{-\frac{1}{2}}\qquad then\quad g'(4) \)

OpenStudy (anonymous):

So 3/2?

OpenStudy (jdoe0001):

yeap, the derivative of a function is the "slope function" for the function if you want to know the value of the slope at "x", just set "x" to that

OpenStudy (jdoe0001):

yes, I got 3/2

OpenStudy (anonymous):

Thank you so much. I really needed to get this right. It's for a takehome quiz and I'm doing terrible in the class. Hopefully this helps. :P Thanks a lot, man.

OpenStudy (jdoe0001):

yw